OpenStudy (anonymous):

Serena drove 40 km on 3 L of gasoline. Which of the following would be an equation to help you find out how far she can travel on a full tank of 50 L? A. 6b.gif B. 6d.gif C. 6a.gif D. 6c.gif

OpenStudy (anonymous):

ok give me sec to et answrs

OpenStudy (anonymous):

(50L/3L)X 40km = 666 + 2/3 km

OpenStudy (anonymous):

ok umm

OpenStudy (anonymous):

B

OpenStudy (anonymous):

o ok ty

OpenStudy (anonymous):

if a car can go 40 km on 3L of gas, then the car should go as many times farther as there are L in the gas tank. 50L/3L = 16 2/3 and 16 2/4 * 40 km equals the ratio in answer B